OpenStudy (anonymous):

Solve using the relate-to-a-simpler-problem strategy The population of Jamesburg is 12,000. Junction City's population is 3/25 of Jamesburg's. What is Junction City's population? A. 1440 B. 3992 C. 4008 D. 100,000

OpenStudy (texaschic101):

I don't know about relating to a simpler problem strategy... but you are wanting to find 3/25 of 12000..." of " means multiply so what is 3/25 * 12000/1 equal ?

OpenStudy (anonymous):

1440?

OpenStudy (texaschic101):

3/25 * 12000/1 = 36000/25 = 1440...yes...you are correct

OpenStudy (anonymous):

Thanks! :D

OpenStudy (texaschic101):

anytime :)